The students need to make 504 muffins,using muffin trays that hold 6 muffins per tray.They plan to put 4 trays at a time into a oven . Each oven takes 20 minutes to bake. How long will they take to make 504 muffins if they will be using 4 ovens


1
Expert's answer
2021-06-22T05:05:24-0400

Number of trays: "N=\frac{504}{6}=84."

In 4 ovens 4*4=16 trays can be putted.

Number of loading: "n=\frac{84}{16}=5.25=6."

Time to bake: "t=20*6=120 min =2 hrs."